OpenStudy (anonymous):

Solve each system using subtraction. y = x + 4 y = 3x Would it be 3x = x + 4? I have that written down, but then I don't know what to do after that.

OpenStudy (anonymous):

so you're solving for x & y?

OpenStudy (anonymous):

Yea, it would be 3x=x + 4. After that, solve for x! Do you understand how you got to 3x = x + 4?

OpenStudy (anonymous):

3x=x+4 3x-x=4 2x=4 x=2 Sub x=2 in y=3x y=3(2) y=6 so x=2 and y=6